OpenStudy (anonymous):

Linear Equations: Kinda need help understanding which formula belongs to which line. Question in the comments.

OpenStudy (anonymous):

\[[Slope = \frac{ a }{ 0 }; x = some.number] <- Horizontal.Line?\]\[[Slope = \frac{ 0 }{ a }; y = some.number] <- Vertical.Line?\]

OpenStudy (inkyvoyd):

think of it this way... if x=3 then that means anywher eon the line x must equal three. thus you can change the y, but the x cannot change because it equals 3. if you can change y, than that means it's vertical

OpenStudy (inkyvoyd):

another way to see it is that slope a/0 is undefined, or +- infinity... that corresponds to a vertical line

OpenStudy (anonymous):

Wait, so [a/0; x = some number] is a horizontal line?

Directrix (directrix):

What does a/0 equal?

OpenStudy (anonymous):

Okay I got it.

Directrix (directrix):

As in 6/0, 10/0 ?

Directrix (directrix):

So, now, what are you saying the answer is?

OpenStudy (inkyvoyd):

no @Sadworld

Directrix (directrix):

>>>Wait, so [a/0; x = some number] is a horizontal line?

Directrix (directrix):

@Sadworld This is not true: >>>Wait, so [a/0; x = some number] is a horizontal line?

OpenStudy (anonymous):

Wait, imma post a pic.

Directrix (directrix):

Good.

OpenStudy (anonymous):

Okay, I reread my material, I got it now. I got it backwards.

Directrix (directrix):

I will share with you what my teacher told me. This vertical and horizontal business is confusing. " So, remember this: Hor-Rye is a trigger "word" to remember that a horizontal line has equation y = a number. "Rye" rhymes with y and is the "trigger" for y =. Ver-Tex is a trigger "word" to remember that a vertical line has equation x = a number. (the x in Tex is the "trigger") She also said this: when zero is underfound, it is undefined. That is sorta loose talk but means if you have a fraction like 5/0, then zero is underfound (in the denominator), so the answer (quotient) is undefined. @Sadworld

Directrix (directrix):

Bottom Line: What you posted in the first reply region is backwards.

Directrix (directrix):

Let's practice: Given the points on a line: (5, 7) and (10, 7), are they points on a vertical or horizontal line? @Sadworld